# Towards Automatic Evaluation and Selection of PHI De-identification Models via Multi-Agent Collaboration

### v1

@@ 1 Introduction
Clinical notes—such as discharge summaries, nursing notes, and radiology or ECG reports—capture rich context like clinical reasoning, observations, and social factors that are often missing in structured electronic health records (Seinen et al., 2025; Tayefi et al., 2021; Zhang et al., 2024). Despite their value for research and downstream applications, clinical notes often contain protected health information (PHI), including names, dates, and addresses, and their reuse is strictly regulated under privacy laws such as HIPAA (Moore and Frye, 2019; Cohen and Mello, 2018). To enable safe sharing and analysis, PHI de-identification is typically applied as a preprocessing step, automatically detecting and removing or replacing personal identifiers—for example, transforming “John Smith was admitted on 03/15/2024 at Hospital A” into “[NAME] was admitted on [DATE] at [HOSPITAL]”.
Existing PHI de-identification (De-id) systems have progressed from handcrafted rules and feature-based machine learning to deep learning and transformer-based models (Kovačević et al., 2024). More recently, large language models (LLMs) have opened new avenues for clinical NLP tasks such as named entity recognition, relation extraction, and note summarization (Lehman et al., 2023; Wu et al., 2024). Studies show that multiple kinds of LLMs can effectively extract PHI entities, generate structured representations from free text, and assist in clinical decision-making (Wu et al., 2025; Liu et al., 2023)
Although LLM-based PHI de-identification has significantly advanced the field, methods for systematically evaluating and automatically selecting the best de-identification model remain underexplored. Conventional evaluation relies on costly, small, institution-specific expert annotations, which impede the creation of reliable, generalizable benchmarks (Altalla’ et al., 2025). LLMs have shown broad prior knowledge and strong zero-shot capabilities (Chang et al., 2024), offering a promising route to assess de-identification models without gold labels. However, using LLMs as verifiers raises new challenges: how to design efficient, calibrated, and robust evaluation pipelines that are insensitive to prompt phrasing or choice of verifier, and how to ensure their judgments are trustworthy across use cases (Chang et al., 2024; de Oliveira et al., 2025).
@@ 2 Related Work
De-identification of PHI has evolved from rule-based systems with handcrafted dictionaries and regexes (Uzuner et al., 2007; Meystre et al., 2010) to statistical learners (CRFs, SVMs) (He et al., 2015; Jiang et al., 2017), then to deep models (BiLSTM-CRF) (Dernoncourt et al., 2017; Tang et al., 2020) and transformer-based architectures (e.g., BioBERT) (Lee et al., 2020; Johnson et al., 2020), improving accuracy while reducing manual feature engineering. Recently, LLMs have been leveraged for many clinical tasks (Xie et al., 2024; Wu et al., 2024), including PHI de-identification, because of their strong contextual reasoning and zero-/few-shot learning abilities (Bhasuran et al., 2025). Recent work has demonstrated the promise of large language models (LLMs) for PHI De-id. For example, DeID-GPT leverages GPT-4 for zero-shot de-identification of medical text, achieving competitive performance without task-specific fine-tuning (Liu et al., 2023). Similarly, Wu et al. (2025) proposed the LPPA framework, which combines synthetic data generation with instruction tuning to fine-tune local LLMs for PHI extraction. LPPA offers a privacy-preserving solution while notably improving recall—a critical metric for safeguarding sensitive health data—and its hybrid models (LPPA-4K and LPPA-5K) are also adopted in our implementation of De-id models. Despite these advances, existing systems still lack robust methods for systematically evaluating and automatically selecting the best De-id model, motivating the development of our proposed framework.
@@
Traditionally, PHI de-identification systems are evaluated against gold-standard corpora such as the i2b2 shared tasks (Stubbs et al., 2015). Metrics typically include precision, recall, and F1-score. However, such datasets are costly to create and limited in scope. Alternative evaluation approaches include weak supervision, ensemble frameworks, and extensible benchmarking systems designed to assess de-identification tools across multiple corpora and settings (Heider and Meystre, 2024). Despite recent progress, robust evaluation without gold-standard annotations, as in this study, is still largely unexplored.
@@
Some researchers has begun to examine the use of LLMs as evaluators, or “judges,” in a range of NLP tasks including summarization, dialogue, and machine translation. These studies show that LLMs can approximate human judgments with promising consistency (Zhu et al., 2023; Hada et al., 2023; Xie et al., 2025; Pan et al., 2024). Specialized evaluator models such as Prometheus incorporate rubric-based scoring to align more closely with human ratings (Kim et al., 2023). However, concerns of bias, calibration, and fairness remain when deploying LLM-based judges at scale (Wang et al., 2023). Our work extends this line of research to the sensitive domain of PHI de-identification, systematically evaluating whether Evaluation Agents can replicate gold-standard and human assessments when ground truth is limited or masked.
@@ 3.3 De-id Models
We employ a diverse set of De-id models—eight LLMs, including two LPPA hybrids trained on synthetic clinical notes (Wu et al., 2025). Each model receives the same prompt and outputs predictions as a JSON dictionary of PHI entities, ensuring a unified, machine-readable format across heterogeneous architectures. Although these models differ in scale, training data, and internal reasoning, our framework treats them purely as black boxes and focuses on rigorous and reproducible evaluation and selection through multi-agent collaboration. The process are described in Eq.(3).
@@ 4.2 LLM Choices
Eight LLMs act as De-id models, including commercial APIs and open-source models, as well as two hybrid LPPA models trained on synthetic clinical notes (Wu et al., 2025). Six LLMs are employed as Evaluation Agents. To ensure fair comparison, all De-id models receive the same prompt and must output a structured JSON dictionary of predicted PHI entities, and Evaluation Agents judge these outputs and return the number of correct predictions in a fixed JSON format.

### v2

@@ 1 Introduction
Clinical notes—such as discharge summaries, nursing notes, and radiology or ECG reports—capture rich context like clinical reasoning, observations, and social factors that are often missing in structured electronic health records (seinen2025using; tayefi2021challenges; zhang2024tacco). Despite their value for research and downstream applications, clinical notes often contain protected health information (PHI), including names, dates, and addresses, and their reuse is strictly regulated under privacy laws such as HIPAA (moore2019review; cohen2018hipaa). To enable safe sharing and analysis, PHI de-identification is typically applied as a preprocessing step, automatically detecting and removing or replacing personal identifiers—for example, transforming “John Smith was admitted on 03/15/2024 at Hospital A” into “[NAME] was admitted on [DATE] at [HOSPITAL]”.
Existing PHI de-identification (De-id) systems have progressed from handcrafted rules and feature-based machine learning to deep learning and transformer-based models (kovavcevic2024identification). More recently, large language models (LLMs) have opened new avenues for clinical NLP tasks such as named entity recognition, relation extraction, and note summarization (lehman2023we; wu2024ontology). Studies show that multiple kinds of LLMs can effectively extract PHI entities, generate structured representations from free text, and assist in clinical decision-making (wu2025large; liu2023deid)
Although LLM-based PHI de-identification has significantly advanced the field, methods for systematically evaluating and automatically selecting the best de-identification model remain underexplored. Conventional evaluation relies on costly, small, institution-specific expert annotations, which impede the creation of reliable, generalizable benchmarks (altalla2025evaluating). LLMs have shown broad prior knowledge and strong zero-shot capabilities (chang2024survey), offering a promising route to assess de-identification models without gold labels. However, using LLMs as verifiers raises new challenges: how to design efficient, calibrated, and robust evaluation pipelines that are insensitive to prompt phrasing or choice of verifier, and how to ensure their judgments are trustworthy across use cases (chang2024survey; de2025study).
@@ 2 Related Work
De-identification of PHI has evolved from rule-based systems with handcrafted dictionaries and regexes (uzuner2007evaluating; meystre2010automatic) to statistical learners (CRFs, SVMs) (he2015crfs; jiang2017identification), then to deep models (BiLSTM-CRF) (dernoncourt2017identification; tang2020identification) and transformer-based architectures (e.g., BioBERT) (lee2020biobert; johnson2020deidentification), improving accuracy while reducing manual feature engineering. Recently, LLMs have been leveraged for many clinical tasks (xie2024promptlink; wu2024ontology), including PHI de-identification, because of their strong contextual reasoning and zero-/few-shot learning abilities (bhasuran2025preliminary). Recent work has demonstrated the promise of large language models (LLMs) for PHI De-id. For example, DeID-GPT leverages GPT-4 for zero-shot de-identification of medical text, achieving competitive performance without task-specific fine-tuning (liu2023deid). Similarly, wu2025large proposed the LPPA framework, which combines synthetic data generation with instruction tuning to fine-tune local LLMs for PHI extraction. LPPA offers a privacy-preserving solution while notably improving recall—a critical metric for safeguarding sensitive health data—and its hybrid models (LPPA-4K and LPPA-5K) are also adopted in our implementation of De-id models. Despite these advances, existing systems still lack robust methods for systematically evaluating and automatically selecting the best De-id model, motivating the development of our proposed framework.
@@
Traditionally, PHI de-identification systems are evaluated against gold-standard corpora such as the i2b2 shared tasks (stubbs2015automated). Metrics typically include precision, recall, and F1-score. However, such datasets are costly to create and limited in scope. Alternative evaluation approaches include weak supervision, ensemble frameworks, and extensible benchmarking systems designed to assess de-identification tools across multiple corpora and settings (heider2024extensible). Despite recent progress, robust evaluation without gold-standard annotations, as in this study, is still largely unexplored.
@@
Some researchers has begun to examine the use of LLMs as evaluators, or “judges,” in a range of NLP tasks including summarization, dialogue, and machine translation. These studies show that LLMs can approximate human judgments with promising consistency (zhu2023judgelm; hada2023large; xie2025kerap; pan2024graphnarrator). Specialized evaluator models such as Prometheus incorporate rubric-based scoring to align more closely with human ratings (kim2023prometheus). However, concerns of bias, calibration, and fairness remain when deploying LLM-based judges at scale (wang2023large). Our work extends this line of research to the sensitive domain of PHI de-identification, systematically evaluating whether Evaluation Agents can replicate gold-standard and human assessments when ground truth is limited or masked.
@@ 3.3 De-id Models
We employ a diverse set of De-id models—eight LLMs, including two LPPA hybrids trained on synthetic clinical notes (wu2025large). Each model receives the same prompt and outputs predictions as a JSON dictionary of PHI entities, ensuring a unified, machine-readable format across heterogeneous architectures. Although these models differ in scale, training data, and internal reasoning, our framework treats them purely as black boxes and focuses on rigorous and reproducible evaluation and selection through multi-agent collaboration. The process are described in Eq.(3).
@@ 4.2 LLM Choices
Eight LLMs act as De-id models, including commercial APIs and open-source models, as well as two hybrid LPPA models trained on synthetic clinical notes (wu2025large). Six LLMs are employed as Evaluation Agents. To ensure fair comparison, all De-id models receive the same prompt and must output a structured JSON dictionary of predicted PHI entities, and Evaluation Agents judge these outputs and return the number of correct predictions in a fixed JSON format.